The displacement of a particle varies with time according to equation. Y=-4/3t ×t+16t+10. How long does the particle take to come to rest

Grade:11

1 Answers

Sheetal kumari
6 years ago
This question is in someway related to differntiationFirstly differntiate this displacement it comes to be -8/3t +16 then velocity = dx/dt and it is 0 as body comes to rest.then t=-16×3÷ -8=6s hope it helps you